Current Yr 1 Yr Ago 2 Yrs Ago At December 31 Assets Cash Accounts receivable, net Merchandise inventory Prepaid expenses PlanComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Compute the rComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Compute the pComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 2a Required 3 Assuming SimoRequired 1 Required 2 Required 2a Required 3 Compute the dividend yield for each year. (Round your answers to 2 decimal place
Answer #1
1) Return on stock holders equity
Net Income - Preferred dividend / Average stockholders equity = Return on stock holders equity
Current Year $             35,424 - 0 / $                                 2,83,823.50 = 12.48%
1 Year Ago $             30,190 - 0 / $                                 2,44,062.50 = 12.37%
2) Price earning ratio
Market price / Earning per share = Price Earning Ratio
Current Year 30 / 2.18 = 13.76
1 Year Ago 28 / 1.86 = 15.05
Simon has higher market expectations for future growth.
3) Dividend Yield
Dividend per share / Market Price per share = Price Earning Ratio
Current Year 0.26 / 30 = 0.87%
1 Year Ago 0.13 / 28 = 0.46%